OpenStudy (anonymous):

show that ∫tanx from 0 to π/4= 1/2In2

OpenStudy (anonymous):

i know that ∫tanx = ∫sinx/cosx = -In(cosx)

OpenStudy (anonymous):

then from there you get -In cos( π/4)

myininaya (myininaya):

omg my i accidently pressed back button damn it

OpenStudy (anonymous):

ohh that sucks lol

OpenStudy (anonymous):

maybe u can tell me where i went wrong in mine?

myininaya (myininaya):

\[-\ln(\frac{\sqrt{2}}{2})=-\ln(\frac{\sqrt{2}}{\sqrt{4}})=-\ln(\sqrt{\frac{2}{4}})=-\ln(\frac{2}{4})^\frac{1}{2}\] =\[-\frac{1}{2}\ln(\frac{1}{2})=\frac{-1}{2}(\ln1-\ln2)=\frac{-1}{2}(0-\ln2)=\frac{1}{2}\ln2\]

myininaya (myininaya):

you remember that cos(pi/4) is sqrt{2}/2 right?

OpenStudy (anonymous):

lol for some reason i made Sqrt(2)xsqrt(2) =4 silly mistake lol

myininaya (myininaya):

lol

OpenStudy (anonymous):

i all makes sense now thank you:)

myininaya (myininaya):

i dont think i ever become your fan let me look

myininaya (myininaya):

ok im your fan i see you so much lol

OpenStudy (anonymous):

lol thank you